OpenStudy (anonymous):

x+1/12=-5/6 I got -11/12 is this right tonight I am just checking work I have done before turning it in hope this is ok

OpenStudy (lgbasallote):

\[\implies x = -\frac 56 - \frac 12\] \[\implies x = -\frac 56 - \frac{3}{6}\] \[\implies x = -\frac{8}{6}\] \[\implies x= -\frac 43\] so i would say you're not right. sorry

OpenStudy (anonymous):

I Got x=-4/5

OpenStudy (anonymous):

my bad I typed it wrong it should be x+1/12=-5/6

OpenStudy (anonymous):

x=-1/12 - 5/6 = -1/12 - 10/12 =-11/12

OpenStudy (anonymous):

cool thanks so much

OpenStudy (baldymcgee6):

If you ever want to check your work... http://www.wolframalpha.com/input/?i=x%2B1%2F12%3D-5%2F6

OpenStudy (anonymous):

thanks just pulled it up